OpenStudy (anonymous):

Simplify completely quantity 2 x minus 24 over 8

OpenStudy (nastech):

i don't really get your question

OpenStudy (anonymous):

2x minus 24 over means : \[\frac{2x - 8}{24}\]

OpenStudy (anonymous):

Or is it : \[2x - \frac{8}{24}\]

OpenStudy (anonymous):

no its\[2x-24/8\]. the numerator is 2x-24 while the denomenator is 8

OpenStudy (anonymous):

I did not get you yet.. Which one above that I have written there? First one or second one??

OpenStudy (nastech):

blondiee please what you wrote are mean 2 different things

OpenStudy (anonymous):

\[\frac{2x-8}{24} \quad Or \quad 2x - \frac{8}{24} \quad ??\]

OpenStudy (anonymous):

its like the first one you just put but replace 8 and 24 making 24 on top with 8 on bottom

OpenStudy (anonymous):

Oh sorry sorry...

OpenStudy (anonymous):

\[2x - \frac{24}{8}\] Now okay??

OpenStudy (anonymous):

no lo okay its\[\frac{ 2x-24 }{ 8 }\]

OpenStudy (anonymous):

So your question must be 2x minus 24 whole divided by 8 Or like this: (2x minus 24) over 8..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know about Distributive Property??

OpenStudy (nastech):

is it all equal to 0

OpenStudy (anonymous):

There is no need of equal to zero here, if it is not there then also it is fine..

OpenStudy (anonymous):

i know about distributive

OpenStudy (anonymous):

So by using that property can you factor out 2 from the numerator part?

OpenStudy (anonymous):

like making it 2(x-12)?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

Now you have : \[\frac{2(x - 12)}{8}\]

OpenStudy (anonymous):

Can you reduce by cancelling 2 with 8??

OpenStudy (anonymous):

Getting @blondiee98 ?

OpenStudy (anonymous):

How many times of 2 is 8 ??

OpenStudy (nastech):

4

OpenStudy (anonymous):

Ha ha ha ha, why you are answering??

OpenStudy (nastech):

i'm also learning

OpenStudy (anonymous):

Nice joke..

OpenStudy (nastech):

lol

OpenStudy (nastech):

I mean it

OpenStudy (nastech):

is that where the simplification ends?

OpenStudy (anonymous):

so would your final answer b\[\frac{ 2(x-12) }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ x-12 }{ 4 }\]

OpenStudy (nastech):

please is that the final answer

OpenStudy (anonymous):

Yeah you are right @blondiee98

OpenStudy (nastech):

thanks a lot

OpenStudy (anonymous):

thank you sooooo much!! :)